Spisak radova Dr Dragutina Djurovica do 1998.godine

1.Jovanovic M., Djurovic D.:Determination de l’heure en 1961.Corrections de la pendule fondamentale et leurs ecarts. Bull.Astron.Obs.Belgrade,XXV,1,1962.

2.Brkic Z., Djurovic D., Jovanovic M.:Determinations de l’heure en 1962. Corrections de la pendule fondamentale et leurs ecarts. Bull.Astron.Obs. Belgrade, XXVI,1,1967.

3.Brkic Z.,Djurovic D., Jovanovic M., Momcilovic R.:Determination de l’heure en 1963.Corrections de la pendule fondamentale et leurs ecarts.Bull. Astron. Obs. Belgrade XXVI, 1, 1967.

4.Brkic Z., Djurovic D., Jovanovic M., Momcilovic R.:Determination astronomique de l’heure en 1964. Bull. Astron. Obs. Belgrade, XXVI, 1, 1967.

5.Brkic Z., Djurovic D., Simic M.: L’heure demi-definitive (systeme TU2) en 0^{s}.0001 a 1964.. Bull. Astron. Obs. Belgrade, XXVI, 1, 1967.

6.Djurovic D.:Les problemes de la determination et de la conservation de l’heure a Belgrade. Proc. National Conf.Astron. Roumanie, Bucarest, 1968.

7.Brkic Z., Djurovic D., Simic M.: L’heure demi-definitive (systeme TU2) en 0^{s}.0001 a 1965. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

8.Brkic Z., Djurovic D., Simic M.: L’heure demi-definitive (systeme TU2) en 0^{s}.0001 a 1966. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

9.Brkic Z., Djurovic D., Simic M.: L’heure demi-definitive (systeme TU2) en 0^{s}0001 a 1967.Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

10.Djurovic D., Simic M.:L’heure demi-definitive (systeme TU2) en 0^{.}0001 a 1968. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

11.Brkic Z., Djurovic D., Jovanovic M., Momcilovic R.: Determination astronomique de l’heure en 1965.Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

12.Brkic Z., Djurovic D., Jovanovic M.:Determination astronomique de l’heure en 1966. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

13. Brkic Z., Djurovic D., Jovanovic M.:Determination astronomique de l’heure en 1967 a 20^{h} TU (systeme TU2).Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

14.Djurovic D.:Determination astronomique de l’heure en 1968 a 20^{h} (systeme TU2). Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

15.Djurovic D.,Radogostic V.:Application de la methode des paires a la determination astronomique de l’heure. Publ. Dept. Astron. Belgrade, 1, 1969.

16.Djurovic D., Radogostic V.:Les erreurs de l’inclinaison et de l’ azimut aux instruments des passages. Publ. Dept. Astron. Belgrade, 1, 1969.

17.Djurovic D., Brkic Z.: Etude des variations de la collimation de l’instrument des passages BAMBERG No 63131 d’ apres des observations astronomiques. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

18. Djurovic D.:Les erreurs systhematiques de la determination de l’epaisseur du contact d’un micrometre impersonnel. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

19.Djurovic D.:L’examen du niveau a bulle de l’ instrument des passages. Bull. Astron. Obs. Belgrade, XXVII, 2, 1969.

20. Djurovic D.: Odredjivanje rektascenzija zvezda iz posmatranja casovnih sluzbi. Publ. Astron. Obs. Beograd, 16,1970.

21.Djurovic D.:Precision des divers types d’instrument appliques aux services de l’heure. Bull. Astron. Obs. Belgrade, XXVIII, F.2,1970.

22. Djurovic D.:Variations saisonnieres de la rotation de la Terre et les systemes de l’heure en 1968.Bull. Astron. Obs. Belgrade, XXVIII,F.2, 1970.

23.Djurovic D.:Determination des coordonnees du pole instantane d’apres les resultats des services de l’heure.Bull. Astron. Obs. Belgrade, XXVIII, F.2,1970.

24.Djurovic D.:Relations entre l’heure atomique,des ephemerides et universelle en 1968. Publ. Dept. Astron. Belgrade, 2, 1971.

25. Djurovic D., Melchior P.:Sur la recherche des termes de maree dans la rotation de la Terre. Bull. classe sci. Accad. Roy. de Belgique,L̄VIII, 1248-1257, 1972.

26.Djurovic D., Radogostic V.: Pritcshiny osnovnih oshibok v astronomitcsheskom opredelenii vremeni. Troudy 18-oj astrometritcsheskoj konferencii SSSR, Poulkovo ,Leningrad, 1973.

27. Djurovic D.:Sur un terme harmonique de 122 jours dans la rotation de la Terre et dans le mouvement du pole. Ciel et Terre, 90, 1, 1974, Bruxelles.

28a.Djurovic D.:Promene brzine Zemljine rotacije. Glas SANU, CCXCI, 37, 1974.

29. Djurovic D.:Determination du nombre de Love k et du facteur Λ= 1+k-l affectants des observations du temps universel. Astron. Astrophys.,47, 1974.

30a. Djurovic D.:Promene longituda i latituda.Relativno pomeranje kontinenata. Glas SANU, C̄CXCI,41, 1975.

31.Djurovic D.:Les termes de maree dans le temps universel.Marees Terrestres: Bull.Obs., Centre international des marees terrestres,IV, 3, Bruxelles, 1975.

32. Djurovic D.:Coordonnees du pole et TU1-TUC pour l’intervalle 1967-1974. Marees Terrestres:Bull. Obs., Centre international des marees terrestres,IV, 3, Bruxelles, 1975.

33. Djurovic D.:Les termes harmoniques dans le mouvement du pole. Bull. Astron. Belgrade, 126,1975.

28b.Djurovic D.:Variations de la vitesse de rotation de la Terre. Bull. SANU, LV, 9, 1976.

30b. Djurovic D.:Variations des longitudes et des latitudes. Deplacement relatif des continents.Bull. SANU, LV,9, 1976.

34. Djurovic D.:Corrections des ascensions droites 245 etoilles du catalogue FK4. Bull. Astron. Belgrade, 127, 1976.

35.Djurovic D.:Amplitudes des composantes de ΔT s et leur depandance de la position geographique.Ciel et Terre, 93, 2, 1977, Bruxelles.

36.Djurovic D.:Sur la conservation d’un systeme de reference dans lequel on exprime x, y et TU1-TUC.Publ. Dept. Astron. Belgrade, 8, 1978.

37. Djurovic D.:Sur l’application de la methode de Nemiro pour le calcul des ascensions droites. Publ. Dept. Astron. Belgrade, 8, 1978.

38.Djurovic D.:Polar coordinates and UT1-UTC from PZT observations. ”Time and Earth’s Rotation”, Proc.IAU. Symp. held in Cadiz (Spain), 1978.

39a. Olevic D., Djurovic D.:Some results of statistical study of Jaccobi constant. Bull SANU, LXIV, 10, 1979.

39b. Olevic D. Djurovic D.:Neki rezultati statistickog ispitivanja Jakobijeve konstante. Glas SANU, CCCXI, 44, 1979.

40. Djurovic D.:Secular and periodical variations of UT2-TAI.Publ. Dept. Astron. Belgrade, 9, 1979.

41.Djurovic D.:Variations of the seasonal inequalities and irregular fluctuations of UT2-TAI.Publ. Dept. Astron. Belgrade, 9, 1979.

42.Olevic D., Djurovic D.: Distributions of the Jaccobi constant and kinetic moment of the Palomar-Leiden Survey asteroids. Publ. Dept. Astron. Belgrade, 9 1979.

43. Olevic D., Djurovic D.:The periodicity of residuals Δρ and Δθ of the double star ADS 11520. Publ. Dept. Astron. Belgrade, 11 , 1981.

44. Djurovic D.:Solar Activity and Earth’s Rotation. Astron. Astrophys.100, 1981.

45. Djurovic D.:Correlation between solar activity and universal time variations. Proc.VI Europ. Meet. ”Sun and Planetary System”, held in Dubrovnik, 1981.

46. Olevic D., Djurovic D.: Contribution to the statistical analysis of Kirkwood’s gapes. Bull. SANU, LXXIX, 12, 1982.

47.Djurovic D.:Preliminary investigation of Belgrade longitude variations. Publ. Dept.Astron. Belgrade, 12, 1983.

48.Djurovic D.:Short-period geomagnetic, atmospheric and Earth’s rotation variations.Astron.Astrophys. 118, 1983.

49.Djurovic D.,Stajic D.:Common short-period cyclic variations of the atmospheric angular momentum and length of day. Bull. SANU 13, 1983.

50.Djurovic D., Stajic D.:Secular and decade fluctuations of the geomagnetic field and universal time UT1. Bull. Astron. Inst. Czechoslovakia 36, 1985.

51.Djurovic D.,Stajic D.: The 75-day cyclic variation of the solar activity and geommagnetic field intensity.Bull. SANU 15, 1986.

52.Djurovic D.: The atmospheric circulation, the Earth’s rotation and solar activity. ”Earth’s Rotation: solved and unsolved problems” , Proc. of NATO Workshop held in Bonas, Toulouse, 1985.

53. Pâquet P., Djurovic D., Techy C.: Agreement in polar motion measurements during MERIT campaign. Proc. of MERIT Workshop held in Columbus, Ohio, U.S.A., 1985.

54.Djurovic D.: The determination of the longitude de Belgrade for the period 1964– 1984. Bull. Astron. Obs. Belgrade 136, 1985.

55.Djurovic D., Stajic D.: Short-period cyclic variations of the solar activity and their geophysical consequences. Publ. Dept. Astron. Belgrade, 13, 1985.

56.Djurovic D.: The accuracy of the chain method solutions. Publ. Dept. Astron. Belgrade, 14, 1986.

57.Djurovic D.: Solar activity, atmospheric circulation and Earth’s rotation. ”Figure and Dynamics of the Earth, Moon and Planets”, Proc. of the IAU/COSPAR Symp. held in Prague, Czechoslovakia, 1986.

58.Djurovic D., Pâquet P.: Short-term variations of the Earth’s rotation and solar activity. ”Earth’s Rotation and Reference Frames in Geodesy and Geophysics”, Proc. of the IAU Symp. held in Washington, U.S.A., 1986.

59.Djurovic D., Pâquet P.: Solar origin of the 50–day fluctuation of the Earth’s rotation and the atmospheric circulation. Astron. Astrophys. 204,1988.

60.Djurovic D., Pâquet P.: A 120-day Oscillation in solar activity and the geophysical phenomena. Astron. Astrophys. 218, 1989.

61.Pâquet P.,Dehant V., Djurovic D.: The 4-month fluctuations observed in the Doppler geodetic height and other solar and geophysical phenomena. Proc. Internat. Geod. Symp. on Satellite Geodetic Positioning held in Las Cruses, U.S.A. , 1989.

62. Djurovic D.:Pseudo-harmonic fluctuations of a random process generated by Labrouste transforms of class Π(s)L. Publ. Dept. Astron. Belgrade, 17, 1989.

63. Djurovic D.: Pseudo-harmonic signal variations generated by Labrouste transforms of the class Π(T). Publ. Dept. Astron. Belgrade 17,1989.

64.Djurovic D., Pâquet P.: A quasi-biennial oscillation in solar activity and geophysical phenomena:preliminary results. Publ. Dept. Astron. Belgrade, 18,1990.

65. Djurovic D.:Solar activity and relationships between astronomy and geosciences.Publ. Dept. Astron. Belgrade, 18, 1990.

66. Djurovic D., Pâquet P.: Variations common to the interplanetary magnetic field,the zonal atmospheric circulation and the Earth’s rotation. Quater.J.Royal Met. Soc.,London, 117, 1991.

67. Djurovic D.: Discrete Fourier transforms of the series with randomly missed data:experimental tests. Bull. Astron. Obs. Belgrade, 145, 1992.

68. Djurovic D.: The least-squares and Wittaker-Robinson-Vondrak method of filter design in the complex deconvolution of data series. Marees Terrestres: Bull. d’Informations,Centre International des Marees Terrestres, Bruxelles, 113, 1992.

69. Djurovic D., Pâquet P.: Quasi-biennial oscillation in green corona activity and Earth’s rotation. Astron. Astrophys. 277, 1993.

70. Djurovic D., Pâquet P.: The frequency composition of the Madden–Julian oscillation. An indication of the solar activity contribution to the global atmospheric circulation and Earth’s rotation. Solar Physics 152, 1994.

71. Djurovic D., Pâquet P., Billiaud A.: New indications for the solar origin of the 50-day cycle in the atmospheric circulation and Earth’s rotation. Astron. Astrophys. 288, 1994.

72. Zhen Nian Gu, Djurovic D, Pâquet P.: The 50-day oscillation in the Earth’s rotation,atmospheric angular momentum and solar activity. Bull. Astron. Obs. Belgrade 152, 1995.

73. Djurovic D.: Present status of Astrometry. Bull. Astron. Obs. Belgrade, 155, 1996.

74. Djurovic D., Pâquet P.: The common oscillations of solar activity, the geomagnetic field and Earth’s rotation. Solar Physics 167, 1996.

75. Damljanovic G., Pejovic N., Djurovic D.: The Chandler nutation for the period 1949–1985. Bull. Astron. Obs. Belgrade 156, 1997.